PLEASE WRITE OUT WHAT FORMULAS YOU USE

7. a. Given the cash flows in actual dollars provided in the following table, convert the cash flows to equivalent cash flows in constant dollars if the base year is time 0. Assume that the market interest rate is 16% and the general inflation rate is at 4% each year. Cash Flow (in Actual Dollars) $20,500 $41,500 $36,500 $55,500 4 Calculate the present worth value of both (actual and constant $) cashflows within the above table. Comment on the obtained results
Answer #1

year                    Cash flow (actual dollars)                      cash flow (constant dollar)
0                      $20,500                                                          $20,500
4                      $41,500                                                   $41,500/1+.04)^4 =   $35474.4
5                      $36,500                                                   $36,500/(1+.04)^5=   $30000.3
7                      $55,500                                                   $55,500/(1+.04)^7=   $42175.4